OpenStudy (fernandez555):

HELP PLEASE!!! keep getting wrong on quiz. if f(x)= x/2 - 2 and g(x) =2x^2 +x-3, find (f+g)(x) choices: a. 2x^2 +3/2x+1 b. 2x^2 +3/2x-5 c. 2x^2 -x/2+1 d. x^2-6 medal!!!!!!!! please include explanation if possible.

OpenStudy (mathmale):

Re-write your " f(x)= x/2 - 2 and g(x) =2x^2 +x-3" as f(x)= x/2 - 2 +g(x) =2x^2 +x-3 --------------------- You should get: (f+g)(x)= ?? combine similar powers of x. For example, x/2 + x = (3/2) x.

OpenStudy (bmk614):

\[\frac{ x }{ 2 }-2+2x^2+x-3\]\[2x^2+\frac{1}{2}x-1x-2-3\] \[2x^2-\frac{ 1 }{ 2 }x-5\]

OpenStudy (fernandez555):

2x^2 + 3x/2 -5 i see where you get the -5, but 1/2x??

OpenStudy (fernandez555):

oh i see, x/2 + x= (3/2)x , i just saw that

OpenStudy (mathmale):

Satisfied? If you are, and have finished this problem, would you please close it. Thx.